the sum of first n term of an A.P is given as n2 -3n.write down the 4th term and the
nth term

Answer by ikleyn(52803)

.

The 4-th term = sum of the first 4 terms  MINUS  sum of the first 3 terms = 


              = (4^2-3*4) - (3^2-3*3) = (16-12) - (9-9) = 4 - 0 = 4.    ANSWER



The n-th term = sum of the first "n" terms  MINUS  sum of the first (n-1) terms = 


             = (n^2-3*n) - ((n-1)^2-3*(n-1)).


Simplify to get the answer.